# Exploring specialization and sensitivity of convolutional neural networks in the context of simultaneous image augmentations

### v1

@@ Appendix D Detailed single-class sensitivity analysis
We propose the single-class sensitivity analysis based on the sensitivity levels of classifying layers (Tables 2, 3). According to the highest prediction offsets generated by specific masking (Tables 4, S7.6), the following augmentations were selected for further analysis: colour adjustments (hue, contrast, saturation, brightness, grayscale), local area distortions (erasing, elliptic local blur), global area distortion (sharpness adjustment) and motions (horizontal flip, rotation). For classes sensible to the motion augmentations, it is natural to expect the presence of symmetry, either vertical line one (horizontal flip) or spherical (rotation). In the case of local area distortions, the sensible classes are expected to have important details and/or elliptic forms in the frame centre (elliptic local blur) along with filled rectangular areas (erasing). As a global area transform, sharping eliminates blurring, emphasizing the contours and shapes, potentially the most important features for predicting classes sensible in that sense. While analyzing the class matching with colour adjustments, it is crucial to recognize the way they affect the input, which may be visualized by changes in colour histograms (channel-wise distributions of intensity values). By adjusting the brightness one is moving the histogram as a whole towards a left or right edge. The hue adjustment shifts histograms against each other, changing the key colour of the whole image. The contrast adjustment increases (decreases) their variances, highlighting (suppressing) the borders between areas of different colours. The saturation parameter controls the image colourfulness by pulling the histograms apart or pushing them together. The limit case for the latter is a grayscale transform where all histograms coincide in the shared area.
@@
Let us start our analysis with the motion augmentations. Masking-based prediction bias supports the single-class sensitivity to a rotation for VGG11 (Shapley, Table S7.6(c)) and ResNet18 (both Shapley and Sobol, Tables S7.6(a),(f)). As expected, classes with object or frame symmetry were encountered (mushroom; hotpot; facepowder; partially paddlewheel; tray and bathtub), but most samples of extracted classes are rotated or contain objects captured from different angles (shoji; abacus; buckle; bathtub; cellular phone), as well as related to the animals photographed in various poses (coyote; hyena; weasel; platypus).
@@
For the horizontal flip, the following common sensitive classes by means of Sobol indices were top-scored for two networks, AlexNet (Table S7.6(a)) and VGG11 (Table S7.6(a),(b)): snowmobile; shield; missile. For the latter two classes, the vertical elongated pattern can be traced. The cliff dwelling objects are aligned with the natural landscape, and can hardly be distinguished from their flipped copies. The same reasoning applies to the rest classes (snowmobile; giant schnauzer; moped; Siberian husky), the objects of which are photographed mainly from the side. I.e., the final layer sensitivity is grading classes according to not only inherent properties but also their capability to be transformed by the corresponding transformation.
@@
The next group of augmentations refers to local distortions. The most notable result is the erasing sensitivity for AlexNet (Sobol, Table 4), where the corresponding objects are rectangular screens and monitors (screen; monitor; television) or rectangular-shaped ones (lighter; pedestal). Similar findings are valid for the ResNet18 (Tables 4, S7.6(c)): Sobol indices top-ranked the same screen; monitor; television classes along with rectangular-shaped wardrobe. At the same time, it is hard to consistently interpret the classes highlighted by Shapley values (AlexNet, Table S7.6(d)), as most of them lack filled rectangular patterns. Classes marked as sensitive to elliptic local blur are slightly easier to explain: for AlexNet (Sobol, Table S7.6(a)) - golf cart and sports car objects have a windscreen centred in the frame or a hood glare; cup images in several samples appear containing an elliptic-like region or glaring spots; black and gold garden spider class contains contrast near-elliptical regions which are important for their recognition. As for ResNet18 (Table S7.6(a),(b)), lacewing fly objects have translucent wings of the same shape, cleaver and stingray are also shaped alike and include spots of light. Similar reasoning may be applied to sensitive classes of VGG11 (Table S7.6(a)): barrel, cask - several objects either have a circular hole in the centre or are captured from the top; parking meter - also contain a round keyhole pattern; coyote - in many pictures the animal is filmed from a distance, occupying a small part in the centre, intersecting with the distortion area; cornet - certain images capture glaring silver or golden musical instruments, some of them are facing the bells into the frame. However, it is difficult to follow the same logic to interpret the rest classes, e.g., cliff dwelling, as sensible to this augmentation.
@@
But the VGG11 is also sensible to cliff dwelling in the context of colour adjustments (contrast, saturation, hue; Sobol, Tables 4, S7.6(a),(c)). This class is represented by pictures of cliffs, canyons, and stone dwellings in the highland area, and many of these objects are exposed to direct sunlight. Some of these photos have been taken from shades, some - from the air. Thus, there is a contrast of light and shades and a prevalence of mono-coloured areas (brown tones, occasionally diluted by green grass and blue sky).
@@
Two networks, ResNet18 (Sobol, Tables 4, S7.6(b)) and VGG11 (Sobol, Tables S7.6(a),(c); Shapley values - unsupported by prediction bias), showed the sensitivity of the jigsaw puzzle class to the contrast adjustment. The images of this class are characterized by the bright colours of the puzzle picture and the sharp edges of its pieces, and the contrast tuning influences the visibility of these edges. Similar logic may explain that walking stick; red-breasted merganser; pay-phone; chain (AlexNet, Sobol), parking meter; gyromitra; waffle iron (VGG11, Shapley), cornet; shovel (VGG11, Sobol), bubble; spotlight; hook, claw (ResNet18, Sobol) are listed as contrast-sensitive classes.
@@
Unlike the previous augmentation, saturation impacts the spacing between channel-wise pixel intensity histograms. The corresponding sensible classes were identified for all networks (Tables 4, S7.6(a),(d),(e)) and share the following basic patterns: there are traces of image processing (e.g., hay; bubble); respective objects contain long colour patches (refrigerator; upright piano) or distinctive monochrome elements (holster; wardrobe); respective objects are set against large monochrome elements of the scene (barbell; sea urchin).
@@
Interestingly, the classes sensible to grayscale transformation differ from the ones for saturation. One is presented in the top-5 list for all three networks, the tree frog class (Sobol, Tables 4, S7.6(a); AlexNet entry is not supported by prediction bias). The corresponding images are in mainly green tones so the inference seems non-relying on the colour feature. The latter characteristic is the same for several other classes of this kind (warthog, scuba diver; Russian wolfhound; - VGG11, Sobol; chain; - ResNet18, Sobol; lakeside - ResNet18, Shapley, Table S7.6(f)). Similarly, the shape seems to be more important for predicting sweatshirt; lifeboat; (ResNet18, Sobol); coral reef; wig (ResNet18, Shapley), and several other classes.
Less obvious is the sensitivity to hue adjustment. There are classes with the typical presence of a diverse environment (often natural) and specifically coloured features, which may be emphasized by shifting the palette: house finch; salamandra (VGG11, Sobol, Table S7.6(a)); bee; cabbage butterfly (ResNet18, Sobol, Table 4). Another recognizable pattern is different shades of characteristic details: digital clock; jellyfish (ResNet18, Sobol). The key colour and palette changes may also confuse other classes with ones characterized by different unnatural colours (e.g., breastplate).
Rich colourful content combined with various lighting conditions is typical for classes sensitive to brightness adjustment: Bedlington terrier; pinwheel; Weimaraner; barrel (VGG11, Sobol; Table S7.6(a)); bookshop; stove; cliff, drop; scabbard (ResNet18, Shapley; Tables S7.6(d),(f)). The objects of classes identified as sensitive to sharpness adjustment are either fine-grained (church; banister; overskirt; AlexNet, Shapley, Table S7.6(d)) or include slim discriminative details (bee eater; leaf beetle; - AlexNet, Shapley; warthog - VGG11, Sobol, Tables S7.6(a),(b)) or their boundary recognition appears to be significant for correct classification (ladle; Alaskian malamute; carton - VGG11, Sobol).
Concluding the findings, single-class sensitivity unsupported by prediction bias seems also correct. That points to the methodology gap which may be filled by incorporating negative feedback, as there are few entries with extremely low values in Tables S7.6.

### v2

@@ Appendix E Detailed single-class sensitivity analysis
We propose the single-class sensitivity analysis based on the class-sensitivity levels of classifying layers (Tables 2–4), with respect to the statistically confirmed outliers listed in the Table 4. SA variables selected for our study can be grouped in the following way: colour adjustments (hue, contrast, saturation, brightness, grayscale), local area distortions (erasing, elliptic local blur), global area distortion (Gaussian blurring, sharpness adjustment) and motions (rolling, horizontal flip, rotation). For classes sensitive to motion augmentations, it is natural to expect the presence of symmetry, either vertical line one (horizontal flip), vertical and horizontal line patterns (rolling), or circular (rotation). In the case of local area distortions, the sensitive classes are expected to have discriminative features such as elliptic forms in the frame centre (elliptic local blur), along with filled rectangular areas (erasing). As a global transform, sharpening eliminates blurring, sometimes introducing a sand texture pattern and emphasizing contours and shapes, potentially the most important features for predicting classes sensitive in that sense, and vice versa for Gaussian blurring. While matching class sensitivity with colour adjustments, it is crucial to recognize how they affect the input, which may be visualized by changes in colour histograms (channel-wise distributions of intensity values). By adjusting the brightness, one is moving the histogram as a whole towards a left or right edge. The hue adjustment shifts histograms against each other, changing the key colour of the whole image. The contrast adjustment increases (decreases) their variances, highlighting (suppressing) the borders between areas of different colours. The saturation parameter controls the image colourfulness by pulling the histograms apart or pushing them together. The limit case for the latter is a grayscale transform where all histograms coincide in the shared area.
Let us start our analysis with the motion augmentations. Several classes identified as sensitive to the rolling transform barely satisfy the expectations (indigo bird, soft-coated wheaten terrier, AlexNet, ILSVRC; mashed potato, chocolate sauce, ResNet18, ILSVRC). A clue to comprehend the matter resides within the subset of samples, which includes the framed pictures, either emerged naturally (e.g., throne room, galley for AlexNet, Sobol, Places365) or artificially generated. The observation is consistent across all networks and datasets, although the pattern is not limited to these selections, indicating that the sensitivity ranking depends on a concrete learning outcome rather than solely on a dataset.
@@
High prediction bias supporting the class sensitivity to rotation was detected for all considered network-dataset pairs. As expected, classes with object or frame circular symmetry were encountered (ILSVRC: mushroom, hotpot, facepowder, partially paddlewheel, tray, and bathtub), but most samples of identified classes are footage from a tilted camera or containing objects captured from different angles (ILSVRC: shoji, abacus, buckle, bathtub, cellular phone; Places365: office, home office), as well as related to the animals photographed in various poses (ILSVRC: coyote, hyena, weasel, platypus).
@@
For the horizontal flip, the following common Sobol-sensitive ILSVRC classes were top-scored for two networks, AlexNet and VGG11: snowmobile, shield, missile. For the latter two classes, the vertical elongated pattern can be traced, as well as for several Places365 classes (basement, staircase, forest road, general store indoor, trench). The cliff dwelling (ILSVRC) objects are aligned with the natural landscape and can hardly be distinguished from their flipped copies. The same reasoning partially applies to the other ILSVRC classes (snowmobile, giant schnauzer, moped, Siberian husky), whose objects are captured mainly from the side. From another point of view, the latter ones are rather asymmetrical, like several sensitive Places365 classes (japanese garden, hospital room, cottage, market outdoor), admitting us to assume that horizontal flip sensitivity is tied to frame symmetry. Since there are also two kinds of individual sensitivity maps of horizontal flip, - sensitive either to the symmetry axis or to the rest field of view, the final-layer sensitivity may be related to both symmetrical and asymmetrical patterns. I.e., the classes are ranked with respect to not only their static inherent properties but also a kind of invariance under the applied transform and its violation.
The next group of augmentations refers to local distortions. The most notable result is the erasing-sensitivity for AlexNet on both datasets (Sobol, Table 5), where the corresponding objects are rectangular screens and monitors (ILSVRC: screen, monitor, television; Places365: home theater, movie theater indoor, television room) or rectangular-shaped ones (lighter, pedestal; phone booth, barndoor). Similar findings are valid for the ResNet18: Sobol indices top-ranked the same screen, monitor, television and home theater, movie theater indoor, television room classes, along with rectangular-shaped wardrobe and barndoor. At the same time, it is hard to consistently interpret the classes highlighted by Shapley values (AlexNet, VGG11), as most of them lack filled rectangular patterns. Probably the more abstract pattern of any large filled monochrome or low-colored areas partially blocking the frame view comes to the forefront. Elliptic local blur sensitivity can be explained by matching the shape and position of specific discriminative details in the frame. For AlexNet, (ILSVRC, Sobol)
@@
golf cart and sports car objects have round lights or wheel often centred in the frame, cup images in several angles appear containing an elliptic-like region, black and gold garden spider class contains contrast near-elliptical regions which are important for their recognition; (Places365, Shapley) - galley and kitchen samples contain rounded metal sinks or kettles, entrance hall - many rounded archs, general store indoor - circled spots of light on the glass bottles. As for ResNet18, (ILSVRC, Sobol): lacewing fly objects have translucent wings of the similar shape, cleaver and stingray are also shaped alike and contain round spots; (Places365, Sobol): lecture room - subset of samples contains microphones with a head of similar shape, youth hostel - lamps with rounded lampshades, greenhouse indoor - rounded plots and planters. Similar reasoning may be applied to sensitive classes of VGG11 (Sobol): barrel, cask - several objects either have a circular hole in the centre or are captured from the top, parking meter - also contain a round filled pattern, coyote - in many pictures the animal is filmed from a distance, occupying a small part in the centre, intersecting with the distortion area, cornet - certain images capture glaring musical instruments facing the bells; (Shapley): scabbard - many short metal daggers, partially resembling the elliptic shape, computer mouse, mosque (the main dome) and wreck - shape similar to the distortion. However, it is difficult to follow the same logic to interpret some other classes, e.g., cliff dwelling, as sensitive to this augmentation.
@@
But the VGG11 is also sensitive to cliff dwelling in the context of colour adjustments (contrast, saturation, hue; Sobol). This class is represented by pictures of cliffs, canyons, and stone dwellings in the highland area, and many of these objects are exposed to direct sunlight. Some of these photos have been taken from shade, some from the air. Thus, there is a contrast of light and shade and a prevalence of mono-coloured areas (brown tones, occasionally diluted by green grass and blue sky).
@@
Two networks, ResNet18 (Sobol) and VGG11 (Sobol and Shapley), showed the sensitivity of the jigsaw puzzle class to the contrast adjustment. The images of this class are characterized by the bright colours of the puzzle picture and the sharp edges of its pieces, and the contrast tuning influences the visibility of these edges. Similar logic may explain that the following classes were listed as contrast-sensitive: walking stick, red-breasted merganser, pay-phone, chain (AlexNet, Sobol, ILSVRC); parking meter, gyromitra, waffle iron (VGG11, Shapley); cornet, shovel (VGG11, Sobol); bubble, spotlight, hook/claw (ResNet18, Sobol, ILSVRC); drumstick, wing, holster (ResNet18, Shapley, ILSVRC). In addition, sensitivity to contrast may be attributed to lighting conditions or color palette affecting the boundary visibility of the key scene objects (AlexNet, Sobol, Places365: television studio, music studio, childs room, amusement arcade, stage indoor; ResNet18, Sobol, Places365: desert vegetation, television studio, kindergarden classroom; AlexNet, Shapley, ILSVRC: napkin, pig, African crocodile).
@@
Unlike the previous augmentation, saturation impacts the spacing between channel-wise pixel intensity histograms. The corresponding sensitive classes were identified for almost all networks (except ResNet18, Places 365, Shapley) and share the following basic patterns: there are traces of image processing visible in several samples (e.g., hay, bubble); respective objects contain long mono-coloured patches (refrigerator, upright piano, screen, water bottle, dough) or distinctive monochrome elements (holster, wardrobe, archaelogical excavation, trench, rotisserie); respective objects are set against large monochrome elements of the scene (barbell, sea urchin); monochrome illumination of the scene or dominating palette of limited colors (discotheque, ice shelf, stage indoor, ice floe, wet bar).
@@
Interestingly, the classes sensitive to grayscale transformation differ from the ones for saturation. One is presented in the top-5 list for all three networks, the tree frog class (Sobol, ILSVRC). The corresponding images are mainly in green tones, so the inference seems non-reliant on the colour feature. The latter characteristic is essential for several other similar classes: warthog, scuba diver, Russian wolfhound - VGG11, Sobol; chain - ResNet18, Sobol, ILSVRC; lakeside - ResNet18, Shapley, ILSVRC; mountain path, lawn, sandbox - AlexNet, Shapley, Places 365. Similarly, the shape seems to be more important for predicting sweatshirt, lifeboat (ResNet18, Sobol, ILSVRC); coral reef, wig (ResNet18, Shapley, ILSVRC), and several other classes. Another possible factor is the presence of grayscale and sepia images among samples (hospital room, archive, music studio, coffee shop, physics laboratory - AlexNet, Sobol, Places 365).
As for hue-sensitive classes, two major patterns related to the key-colour and palette changes can be observed. The samples of the first kind contain the dominant colour of the background with shade varying - including the variability caused by artificial illumination or differences in the natural light level (e.g., agaric, missile, Chihuahua, house finch, salamandra, breastplate, bee - ILSVRC; stage indoor, rainforest, ruin, beer hall, plaza, discotheque - Places365). The second recognizable pattern emphasizes the colour of the key object or its critical feature, especially for close-up shots (digital clock; botanical garden, vegetable garden, florist shop indoor). Several hue-sensitive classes combine both patterns (slum, jeep, jellyfish; recreation room).
Similarly, one may isolate the following pair of typical patterns for brightness-sensitive classes: rich colorful content (grey wolf, howler monkey, bookshop - ILSVRC; childs room, playroom, candy store, art studio, pizzeria, bazaar indoor - Places365); various luminosity levels, often lighting-dependent (worm fence, boxer, cliff/drop; creek, patio, crosswalk, rock arch, corral); their combination (pinwheel, plate, king snake; general store outdoor).
The latter group of augmentations consists of Gaussian blur and sharpness adjustment, both impacting the same high-level image feature, sharpness. Gaussian blur cuts off fine details, complicating the class discrimination. Such fine-detailed classes sensitive to blurring are: throne room, temple asia, server room (Places365). Similarly to them, several other classes include low resolution samples (broom, hair slide - ILSVRC; restaurant, market indoor, children room, shoe shop, kitchen, coffee shop, playroom - Places365). Others contain defocused or blurred samples (scuba diver, bubble, green snake - ILSVRC). Finally, the sensitivity may be due to the inherent coarse properties of the objects (mask, coho salmon, flat-coated retriever, affenpinscher, dingo; shed, barndoor, driveway, basement, fabric store).
Sharpness adjustment may introduce a sandy texture or amplify grain artifacts. Several classes identified as sensitive to this augmentation possess similar patterns (warthog, coho salmon, Alaskan malamute, poncho, harvester/reaper, face powder; excavation, lawn). There are other classes with fine details, as for blurring augmentation, and images of these classes are frequently diverse, colourful, with numerous details, and are prone to being confused with other classes where similar details emerged due to the augmentation artifacts (church; delicatessen, bazaar outdoor, church outdoor, temple asia). Other sensitive classes feature various ripples and edges which the transformation emphasizes (banister, bee eater, leaf beetle, overskirt, carton, Windsor tie, grasshopper; harbor, elevator lobby).